### v2

@@ 3 Joint Liability Contracts under a Mean–Variance Utility Framework
3.1 The optimal $ESE$ score with different levels of risk aversion
When borrowers consider potential projects, they would like to know how likely it is that they will be able to earn a high return or experience a substantial loss. The variability of profits, or the range of possible returns, is an important consideration for these borrowers. If the potential variability of returns is too high or the risk of substantial losses is too large, the borrower may decide not to invest in the venture. To the best of our knowledge, the current literature on joint liability models focuses only on the expected return or expected profit. This means that they concentrate solely on the central point of an individual’s profit without taking into account the distribution of potential profits or the variability of these returns. However, focusing only on the expected return can be limiting, as it does not account for the risks. In this paper we incorporate the risk (proxied by variance) associated with expected profit into our utility function and conduct further analysis of mean–variance utility in the context of the individual ESE–joint liability model. Specifically the mean–variance utility function for an individual farmer can be incorporated in his/her ex–ante expected profit function in Equation (3.1), where $P$ (a random variable) represents the expected profit and $\gamma$ represents a risk aversion parameter of the farmer
@@ 3.1 The optimal $ESE$ score with different levels of risk aversion
-- $\frac{\gamma k}{2}\left\{\left(2e-4e^{3}\right)(p\overline{Y}-w)^{2}+\left(1-4e+6e^{2}-4e^{3}\right)(p\overline{Y}+p\underline{Y}-2w)^{2}\right.$
$\left.-2\left(3e^{2}-4e^{3}\right)(p\overline{Y}-w)(p\overline{Y}+p\underline{Y}-2w)\right\}$ (3.2)
$=$ 00 0
Unfortunately the resulting analytical solution of the ESS score involves a rather complex expression and is difficult to interpret intuitively; so we resort to simulation studies to reveal several insightful patterns. Recall that the probability of achieving a success in a project is given by $e=k\times ESE+b$, wherein the parameters $k$ and $b$ are predetermined. Subsequently the probability may be disaggregated into two constituent components. Specifically the product of $k$ and $ESE$ represents a contribution of a personal impact towards a project success, whereas $b$ represents a constant contribution of exogenous factors beyond the individual control. The value of $k$ is utilized to calibrate the impact of $ESE$ on a project success, while $b$ pertains to the influence of exogenous factors, including climatic and geographical conditions. To determine the value of $b$, an Actuaries Climate Index (ACI) for each region can in principle be utilized as a reference. Briefly the ACI is based on an analysis of six climate indicators, including high and low temperatures, heavy precipitation, drought, and sea level rise. The index measures changes in these indicators over time and across regions, providing indication on how the climate is changing in different parts of a region. It is commonly understood that favorable climate conditions have a positive effect on crop yields, which, in turn, increases the probability of project success. Conversely, unfavorable climate conditions have an adverse impact on the likelihood of success. Hence, we delve deeper into the impact of varying climate conditions caused by climate change on ESE score requirements. By changing the farmer’s degree of risk aversion from 0 to 1 under different cost parameters, a series of simulations are conducted for a high success probability project under favorable climate conditions (where $b=0.7$ in Figure 3), a neutral probability project under neutral climate conditions (where $b=0.5$ in Figure 4), and a relatively low probability project under unfavorable climate conditions (where $b=0.3$ in Figure 5). After observing and comparing these three figures, we conclude our analysis with three findings below. In general any factor that can increase the success probability of the project and the repayment probability can directly or indirectly reduce the requirements of the $ESE$ score. By examining Figures 3, 4, and 5, we notice that, as the probability of success decreases, the $ESE$ score requirement continues to increase if the control cost remains constant; i.e., the probability of a project’s success is inversely proportional to the required ESE score. This is because a higher success rate leads to an increased farmers’ ability to repay their loans, thus lowering the risk of default for the insurance company. As a result the insurance company can lower its $ESE$ score requirements when lending to farmers. From inspecting any one of the three following figures, we see that, as the degree of risk aversion increases, the required ESE score decreases monotonically. This is because an increase in farmers’ risk aversion score leads to a more cautious contract behavior, reducing the probability of a strategic default and increasing the probability of repayment. In other words, a higher degree of risk aversion on the part of the farmers implies a lower level of risk for the insurance company, leading to a lower ESE score requirement. Again, from inspecting any one of the three figures, we can conclude that an increase in the cost parameter will result in a higher requirement for the ESE score. An increase in the cost parameter raises the production expenditure of farmers and burdens their economic resources, thus reducing their profits. Therefore, to avoid a high default probability associated with high costs, the corresponding ESE score requirement should increase. Hence, as the cost parameter increases, the insurance company should raise its ESE score requirement to reduce the risk of default among farmers.
@@
Finally we arrive at the question of how changes in crop yields can impact the required $ESE$ score. To answer this question, we vary the values of yields and observe the effect on the required ESE score. Specifically we consider two scenarios: a ‘Low Yield Case’ where $\overline{Y}=1,000$ and $\underline{Y}=500$, and a ‘High Yield Case’ where $\overline{Y}=600$ and $\underline{Y}=300".$ The results of our simulation in Figure 6 show that there is a negative correlation between crop yields and the required ESE score, i.e., as crop yields increase, the required ESE score decreases. Conversely, as crop yields decrease, the required ESE score increases. When a borrower has lower yields, this can be taken as an indication of a lower income and a potentially higher risk of default; in other words, a higher requirement for borrowers with lower yields is a reflection of the lender’s risk management strategy. By setting stricter requirements for the borrowers with lower yields, the lender is attempting to minimize their exposure to financial risk and ensure that the borrowers are able to repay their loans on time and in full.
@@ 4 Conclusions
In this paper we have incorporated a novel concept coined as an individual–ESE score into a joint liability model for the purpose of evaluating the sustainability and creditworthiness of individual farmers acting as borrowers in the model. To achieve this objective we initially considered the expected profit of an individual borrower resulting from the specification of his/her utility function. Our primary objective was to determine the optimal group size that maximizes the benefit to borrowers while ensuring that the lender breaks even. Additionally we sought to derive the optimal individual–ESE score, which can provide several benefits to both the lending companies and the farmers as borrowers, such as a better understanding of the borrower’s risk profile, information about the optimal group size for lending, and an improved access to credits. The inclusion of this score in the credit evaluation processes can lead to better–informed lending decisions and sustainable development outcomes. We listed the main results of this paper conveniently in five propositions. Proposition 1 focused on finding an optimal individual-ESE score. Proposition 2 aimed at collapsing two constraints on the loan ceiling to a single constraint. Proposition 3 demonstrated that a higher individual-ESE score leads to a higher loan ceiling. Proposition 4 revealed that the optimal individual-ESE score and group size $n$ have a negative relationship. Finally Proposition 5 showed that the optimal individual–ESE score converges to a constant level when the group size reaches a large enough number. As a next step our study incorporated the risk associated with an expected profit into the utility function of the farmers and analyzed the resulting mean-variance utility framework in the context of the individual $ESE$–joint liability model. The study additionally disentangled the probability of achieving success in a project into two constituent components, namely the impact of personal factors (represented by $k\times ESE$) and exogenous factors beyond individual farmer’s control (represented by $b$), such as climatic and geographical conditions. Our findings reveal that the probability of a project’s success is inversely proportional to the required ESE score. This implies that a higher likelihood of a project success leads to an increased farmers’ ability to repay their loans, thereby lowering the risk of default for the insurance company. As a result the insurance company can lower its $ESE$ score requirements when lending to high-quality farmers. We also found that an increase in farmers’ risk aversion score leads to a more cautious contract behavior, reducing the probability of a strategic default and increasing the probability of repayment. Hence, as the risk–aversion score increases, the corresponding $ESE$ score requirement decreases. Finally an increase in the cost parameter raises the production expenditure of the farmers and places a burden on their economic resources which reduces their profits. Therefore, to avoid a high default probability associated with high costs, the corresponding requirement of an ESE score should be increased. Subsequently we explored the effects of climate change on the required ESE score in order to ensure a successful project implementation. Our simulation results demonstrated that a higher likelihood of a project success, driven by high yields resulting from favorable climate conditions, leads to a lower $ESE$ score requirement in the proposed individual ESE–joint liability model. This suggests that incorporating climate scenarios into the evaluation of creditworthiness can result in better–informed lending decisions and ultimately contribute to more favorable sustainable development outcomes. As an avenue for future research, this study paves the way for an exploration of a partially joint liability’s legal framework which distributes shared responsibility among parties for specific obligations. Unlike a complete joint liability frameworrk, a partially joint liability framework assigns a proportionate debt share to each party. Moreover the concept of a flexible joint liability introduces aspects of innovation. Participants begin with a collaborative lending similar to a standard joint liability contract but, in strategic default scenarios, opt for a punitive phase rather than immediate maximum penalties. After a defined period ($T$ as previously mentioned), defaulters are reintegrated, which offers a path for forgiveness and adherence to the flexible joint liability contract. The diverse joint liability arrangements highlight the need for a thorough and separate analysis. Last but not least, our review of the literature has revealed that while traditional credit metrics concentrate on financial and economic factors, assessing sustainability requires taking into account environmental and social factors as well. Therefore, as a first step of analysis, we have suggested the concept of an individual-ESE score that integrates ESE considerations to offer a more encompassing appraisal of the farming sector. It follows that a potentially most consequential avenue for a future research is to eventually develop a credit scoring system that addresses the unique requirements of the agricultural sector by applying literature–based metrics for agricultural credit scoring systems to the construction of the scores. By compiling and adjusting these metrics, we hope to eventually be able to create a comprehensive credit scoring system with adaptation to climate change for the agricultural sector that addresses limitations of the ESG scores as they are applied to the agricultural sector. This is an important task to take up but is clearly beyond the scope of the present study.

### v3

@@ 4 Joint Liability Contracts under a Mean–Variance Utility Framework
When borrowers assess potential projects, they would like to know how likely it is that they will be able to earn a high return or experience a substantial loss. The variability of profits, or the range of possible returns, is an important consideration for these borrowers. If the potential variability of returns is too high or the risk of substantial losses is too large, the borrower may decide not to invest in the venture. To the best of our knowledge, the current literature on joint liability models focuses only on the expected return or expected profit. This means that they concentrate solely on the central point of an individual’s profit without taking into account the distribution of potential profits or the variability of these returns. However, focusing only on the expected return can be limiting, as it does not account for the risks. In this paper, we incorporate the risk (proxied by variance) associated with expected profit into our utility function and conduct further analysis of mean-variance utility in the context of the individual ESE–joint liability model. Specifically, the mean-variance utility function for an individual farmer can be incorporated in his/her ex–ante expected profit function in Equation (4.1), where $P$ (a random variable) represents the expected profit and $\gamma$ represents a risk aversion parameter of the farmer. The choice to limit the group size to two in this section is intended only to serve as a stylized example to illustrate the key concepts of the mean-variance utility framework. Extending the analysis to larger group sizes introduces complexity further, as it requires simultaneously accounting for both the return and risk components across multiple group members. While such an extension would provide a more general analysis, it is beyond the scope of this study and is left for future work.
@@
- $\frac{\gamma k}{2}\left\{\left(2e-4e^{3}\right)(p\overline{Y}-w)^{2}+\left(1-4e+6e^{2}-4e^{3}\right)(p\overline{Y}+p\underline{Y}-2w)^{2}\right\}$
- $\frac{\gamma k}{2}\{2\left(3e^{2}-4e^{3}\right)(p\overline{Y}-w)(p\overline{Y}+p\underline{Y}-2w)\}$ (4.2)
$=$ $0$
Unfortunately, the resulting analytical solution of the ESE score involves a rather complex expression and is difficult to interpret intuitively; so we resort to simulation studies to reveal several insightful patterns.
Recall that the probability of achieving success in a project is given by $e=k\times ESE+b$, wherein the parameters $k$ and $b$ are predetermined. Subsequently, the probability may be disaggregated into two constituent components. Specifically, the product of $k$ and ESE represents a contribution of a personal impact towards a project’s success, whereas $b$ represents a constant contribution of exogenous factors beyond individual control. The value of $k$ is utilized to calibrate the impact of ESE on a project’s success, while $b$ pertains to the influence of exogenous factors, including climatic and geographical conditions.
To determine the value of $b$, an Actuaries Climate Index (ACI) for each region can in principle be utilized as a reference. Briefly, the ACI is based on an analysis of six climate indicators, including high and low temperatures, heavy precipitation, drought, and sea level rise. The index measures changes in these indicators over time and across regions, providing an indication of how the climate is changing in different parts of a region.
It is commonly understood that favorable climate conditions have a positive effect on crop yields, which, in turn, increases the probability of project success. Conversely, unfavorable climate conditions have an adverse impact on the likelihood of success. Hence, we delve deeper into the impact of varying climate conditions caused by climate change on ESE score requirements. By changing the farmer’s degree of risk aversion from 0 to 1 under different cost parameters, a series of simulations are conducted for a high success probability project under favorable climate conditions (where $b=0.7$ in Figure 3), a neutral probability project under neutral climate conditions (where $b=0.5$ in Figure 4), and a relatively low probability project under unfavorable climate conditions (where $b=0.3$ in Figure 5). After observing and comparing these three figures, we conclude our analysis with three findings below. In general, any factor that can increase the success probability of the project and the repayment probability can directly or indirectly reduce the requirements of the ESE score.
By examining Figure 3, Figure 4, and Figure 5, we notice that, as the probability of success decreases, the ESE score requirement continues to increase if the control cost remains constant; i.e., the probability of a project’s success is inversely proportional to the required ESE score. This is because a higher success rate leads to an increased farmers’ ability to repay their loans, thus lowering the risk of default for the insurance company. As a result, the insurance company can lower its ESE score requirements when lending to farmers.
From inspecting any one of the three following figures, we see that, as the degree of risk aversion increases, the required ESE score decreases monotonically. This is because an increase in farmers’ risk aversion score leads to a more cautious contract behavior, reducing the probability of a strategic default and increasing the probability of repayment. In other words, a higher degree of risk aversion on the part of the farmers implies a lower level of risk for the insurance company, leading to a lower ESE score requirement.
Again, from inspecting any one of the three figures, we observe that an increase in the cost parameter will result in a higher requirement for the ESE score. An increase in the cost parameter raises the production expenditure of farmers and burdens their economic resources, thus reducing their profits. Therefore, to avoid a high default probability associated with high costs, the corresponding ESE score requirement should increase. Hence, as the cost parameter increases, the insurance company should raise its ESE score requirement to reduce the risk of default among farmers.
@@
Finally, we arrive at the question of how changes in crop yields impact the required ESE score. To answer this question, we vary the values of yields and observe the effect on the required ESE score. Specifically we consider two scenarios: a ‘Low Yield Case’ where $\overline{Y}=1,000$ and $\underline{Y}=500$, and a ‘High Yield Case’ where $\overline{Y}=600$ and $\underline{Y}=300$". The results of our simulation in Figure 6 show that there is a negative correlation between crop yields and the required ESE score, i.e., as crop yields increase, the required ESE score decreases. Conversely, as crop yields decrease, the required ESE score increases. When a borrower has lower yields, this can be taken as an indication of a lower income and a potentially higher risk of default; in other words, a higher requirement for borrowers with lower yields is a reflection of the lender’s risk management strategy. By setting stricter requirements for borrowers with lower yields, the lender is attempting to minimize their exposure to financial risk and ensure that the borrowers are able to repay their loans on time and in full.
@@ 5 Conclusions
In this paper, we have incorporated a novel concept coined as an individual–ESE score into a joint liability model for the purpose of evaluating the sustainability and creditworthiness of individual farmers acting as borrowers in the model in the face of climate change. To achieve this objective we initially considered the expected profit of an individual borrower resulting from the specification of his/her utility function. Our primary goal was to determine the optimal individual–ESE score and to investigate its relationship with key factors such as group size and loan ceiling. By understanding these relationships, we aimed at providing useful insights that benefit both the lending companies and the farmers acting as borrowers. The optimal individual–ESE score serves as a valuable tool for lenders by offering a better understanding of borrowers’ risk profiles and informing decisions about appropriate group sizes for lending. For farmers, it can facilitate improved access to credit, thereby fostering greater financial inclusion. Integrating this score into credit evaluation processes not only supports better informed lending decisions but also promotes sustainable development outcomes by aligning financial practices with environmental and social considerations. We listed the main results of this paper conveniently in five propositions. Proposition 1 focused on finding an optimal individual-ESE score. Proposition 2 aimed at collapsing two constraints on the loan ceiling into a single constraint. Proposition 3 demonstrated that a higher individual-ESE score leads to a higher loan ceiling. Proposition 4 revealed that the optimal individual-ESE score and group size $n$ have a negative relationship. Finally, Proposition 5 showed that the optimal individual–ESE score converges to a constant level when the group size reaches a large enough number.
Our study went on to incorporate the risk associated with an expected profit into the utility function of the farmers and analyzed the resulting mean-variance utility framework in the context of the individual ESE–joint liability model. The study additionally disentangled the probability of achieving success in a project into two constituent components, namely the impact of personal factors (represented by $k\times ESE$) and exogenous factors beyond individual farmer’s control (represented by $b$), such as climatic and geographical conditions. Our findings reveal that the probability of a project’s success is inversely proportional to the required ESE score. This implies that a higher likelihood of a project’s success leads to an increased farmers’ ability to repay their loans, thereby lowering the risk of default for the insurance company. As a result, the insurance company can lower its ESE score requirements when lending to high-quality farmers. We also found that an increase in farmers’ risk aversion score leads to a more cautious contract behavior, reducing the probability of a strategic default and increasing the probability of repayment. Hence, as the risk–aversion score increases, the corresponding ESE score requirement decreases. Finally, an increase in the cost parameter raises the production expenditure of the farmers and places a burden on their economic resources which reduces their profits. Therefore, to avoid a high default probability associated with high costs, the corresponding requirement of an ESE score should be increased.
Lastly, we explored the effects of climate change on the required ESE score in order to ensure a successful project implementation. Our simulation results demonstrated that a higher likelihood of project success, driven by high yields resulting from favorable climate conditions, leads to a lower ESE score requirement in the proposed individual ESE–joint liability model. This suggests that incorporating climate scenarios into the evaluation of creditworthiness can result in better-informed lending decisions and ultimately contribute to more favorable sustainable development outcomes.
As an avenue for future research, this paper paves the way for an exploration of a partially joint liability legal framework that distributes shared responsibility among parties for specific obligations. Unlike a complete joint liability framework, a partially joint liability framework assigns a proportionate debt share to each party. Moreover, the concept of a flexible joint liability introduces aspects of innovation. Participants begin with a collaborative lending similar to a standard joint liability contract but, in strategic default scenarios, opt for a punitive phase rather than immediate maximum penalties. After a defined period ($T$ as previously mentioned), defaulters are reintegrated, which offers a path for forgiveness and adherence to the flexible joint liability contract. The diverse joint liability arrangements highlight the need for a thorough and separate analysis. Future research could explore the application of various joint liability frameworks to enhance understanding and implementation in this area.
Furthermore, to extend the model to appeal to the actuarial domain, it can incorporate key actuarial principles such as risk pooling, premium calculation, and correlated risks. Introducing loss distributions and reinsurance mechanisms would align the model with actuarial practices for assessing extreme losses and managing shared liabilities. The model can also include dynamic economic factors like interest rate and income to capture long-term liability risks, while stochastic optimization techniques could be employed to balance group composition and heterogeneity. Additionally, actuarial approaches for valuing contingent liabilities, setting capital requirements, and integrating ESG factors would make the model more robust and applicable to real-world financial and risk management scenarios. These enhancements would provide practical tools for actuaries to assess, mitigate, and optimize joint liability frameworks effectively.